A charge of 100 elementary charges is equivalent to?
(1) 1.60 x 10-21 C (3) 6.25 x 1016 C
(2) 1.60 x 10-17 C (4) 6.25 x 1020 C
Proof: Explain your reasoning/show calculations.
1
Expert's answer
2020-04-21T10:56:23-0400

Elementary charge is the charge of electron: "e =1.6 \cdot 10^{-19} C".

Hence, the charge of 100 elementary charges is "100 \cdot 1.6 \cdot 10^{-19} C = 1.6 \cdot 10^{-17} C".

Correct answer is 2)